BROADENING, SHIFTS AND LINE MIXING IN THE ^{13}CH_{4} ν_{4} BAND

\underline{D. CHRIS BENNER}, V. MALATHY DEVI, Department of Physics, College of William and Mary, Williamsburg, VA 23187-8795; M.A.H. SMITH, Science Directorate, NASA Langley Research Center, MS 401A, Hampton, VA 23681-2199.

We have reanalyzed our air-broadened spectra of the ν_{4} band of ^{13}CH_{4} by employing a multispectrum nonlinear least squares fitting procedure.^{a} To fit the highest pressure spectrum well, it was sometimes necessary to invoke line mixing. In these cases the off-diagonal relaxation matrix elements were determined directly from the fits. The data contained four room temperature ^{13}CH_{4} spectra together with up to sixty ^{12}CH_{4} spectra including both self- and air broadening taken at different pressures and temperatures. The four ^{13}CH_{4} spectra included one low-pressure spectrum of a 99 percent 13C-enriched CH_{4} in a 0.0985-cm Pyrex absorption cell and three spectra with ∼0.5 percent volume mixing ratio of ^{13}CH_{4} in air contained in a 50-cm cell also made of Pyrex. Air-broadened halfwidth and air-induced pressure shift coefficients were measured for nearly 200 transitions of the ^{13}CH_{4} ν_{4} band with J” as large as 16. Line mixing coefficients were retrieved where appropriate. A comparison of present results with previous measurements will be made. This research was conducted at the College of William and Mary and Langley Research Center under contract with the National Aeronautics and Space Administration

aD. Chris Benner et al., JQSRT \underline{53}, 705-721 (1995).